OpenStudy (anonymous):

integrate (secx)^4 * sqrt (tanx) dx

OpenStudy (anonymous):

I think you have to use Trig sub

OpenStudy (anonymous):

or you might be able to do this

OpenStudy (anonymous):

this is just a test idk if it will work \[\int sec^2 x* sec^2 x*\sqrt{tanx} dx\]

OpenStudy (anonymous):

\[cos^2 x + sin^2 x= 1\] \[1+tan^2 x=sec^2 x\]

OpenStudy (anonymous):

you're correct, i'm close, but not exactly the ansewr

OpenStudy (anonymous):

\[\int sec^2 x * (tan^2 x + 1) * \sqrt{tan x}dx\]

OpenStudy (anonymous):

not trig sub, u-sub

OpenStudy (anonymous):

yeah then let u = tanx; du = (secx)^2

OpenStudy (anonymous):

you could use trig subs

OpenStudy (anonymous):

trig sub will work with any trig integral... it's just a pain for something that can be done easier anyways \[\int sec^2 x *(tan^{3/2} x+ \sqrt{tan^{1/2}x})\]

OpenStudy (anonymous):

\[\int sec^2 x*tan^{3/2}x dx + \int sec^2 x*tan^{1/2}xdx\]

OpenStudy (anonymous):

\[\int sec^2 x* (tanx)^{3/2} dx + \int sec^2 x*(tanx)^{1/2}\]

OpenStudy (anonymous):

let \[u=tanx\] \[du=sec^2 x\] \[\int u^{3/2}du+\int u^{1/2}du\]

OpenStudy (anonymous):

i see where i went wrong

OpenStudy (anonymous):

ok =]

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

∫sec2x∗(tan3/2x+tan1/2x−−−−−−√)

OpenStudy (anonymous):

how does tan^2 (x) * tan^(1/2) x = tan^(3/2) (x)?

OpenStudy (anonymous):

ehh it's supposed to be a 5/2

OpenStudy (anonymous):

so your answer wil lbe 7/2

random231 (random231):

wow